OpenStudy (anonymous):

Find the limit of the sequence An=(9n+1)/sqrt(25n^2-36) And An=n^2/(6^n+8)

OpenStudy (anonymous):

\[\lim_{n\to\infty}A_n=\lim_{n\to\infty}\frac{9n+1}{\sqrt{25n^2-36}}=\sqrt{\lim_{n\to\infty}\frac{81n^2+18n+1}{25n^2-36}}\\\ \ \ \ \ \ \ \ \ \ \ =\sqrt{\lim_{n\to\infty}\frac{162n+18}{50n}}=\sqrt{\lim_{n\to\infty}\frac{162}{50}}=\sqrt{\frac{162}{50}}=\sqrt\frac{81}{25}=\frac95\]

OpenStudy (anonymous):

Thank you! So the 2nd one doesn't have a limit?

OpenStudy (anonymous):

Not sure, I was just too lazy to find it... you can determine that by finding the limit of the ratio between terms and seeing if it's 1 :-)

OpenStudy (anonymous):

Lol thanks

OpenStudy (ammarah):

yup what he did...